Find voltage drop and current across the 2 ohm resistormi 28v++ 20 v
Answer #1

no in 28 V 321 Let Voltage across 2n is v. by KCL, [ussing modal analysis] V-28+ 2 VE 4(-7)=0 V-28+2V + 4 V-28 = 0 70-56 = 0